# EmoAttack: Emotion-to-Image Diffusion Models for Emotional Backdoor Generation

### v1

@@ Abstract
Text-to-image diffusion models can create realistic images based on input texts. Users can describe an object to convey their opinions visually. In this work, we unveil a previously unrecognized and latent risk of using diffusion models to generate images; we utilize emotion in the input texts to introduce negative contents, potentially eliciting unfavorable emotions in users. Emotions play a crucial role in expressing personal opinions in our daily interactions, and the inclusion of maliciously negative content can lead users astray, exacerbating negative emotions. Specifically, we identify the emotion-aware backdoor attack (EmoAttack) that can incorporate malicious negative content triggered by emotional texts during image generation. We formulate such an attack as a diffusion personalization problem to avoid extensive model retraining and propose the EmoBooth. Unlike existing personalization methods, our approach fine-tunes a pre-trained diffusion model by establishing a mapping between a cluster of emotional words and a given reference image containing malicious negative content. To validate the effectiveness of our method, we built a dataset and conducted extensive analysis and discussion about its effectiveness. Given consumers’ widespread use of diffusion models, uncovering this threat is critical for society.
Machine Learning, ICML
@@ 1 Introduction
In recent years, substantial progress has been achieved in the realm of text-to-image diffusion models. These models, which govern generation through textual prompts (Rombach et al., 2022; Saharia et al., 2022), have introduced unparalleled creativity and functionality across various domains, including art generation and media production. Notably, text-to-image methods also empower individuals to articulate descriptive opinions visually. This capability can potentially enhance the modes of interaction between humans and human-machines. One person could entertain himself or interact with other persons by providing their text descriptions and generating images that could be easily understood.
Meanwhile, emotions constitute an essential element of the human experience, influencing various facets of our lives and encompassing human behaviors (Trampe et al., 2015). Humans usually use emotional words in the text descriptions to express their emotions implicitly or explicitly. For example, if a person feels sad and we ask him to describe what he sees, he may use sadness-related words like ‘sorrowful’, ‘heartbroken’, ‘dejected’, etc. Given the importance of emotion within the human description and the progress text-to-image methods, we unveil a latent risk of using diffusion models to generate images, that is, using emotion as a trigger to introduce malicious negative contents, potentially eliciting unfavorable emotions in users, which is an unrecognized risk previously. We denote such a task as the emotion-aware backdoor attack (EmoAttack).
Backdoor attacks against diffusion models have been studied in previous works (Chou et al., 2023a; Zhai et al., 2023; Chou et al., 2023b). (Chou et al., 2023a) studies the backdoor attack against image-to-image tasks using specific images as the trigger. (Chou et al., 2023b) and (Zhai et al., 2023) can set texts as triggers. However, these backdoor attacks consider subject-related discrete words (e.g., ’cat,’ ’dog’) as triggers, which diverge significantly from the prerequisites of EmoAttack and are not adaptable to its requirements since emotional triggers tend to be more abstract and nuanced. For example, a wide range of words can represent the emotion ‘sadness’ (See Fig. 1).
To overcome this challenge, we introduce EmoBooth. Specifically, we present a method to represent emotions for a given subject through diverse sentences generated by ChatGPT in a latent space. We then create backdoor texts by decoding samples around the clustering center. Furthermore, we suggest an emotion injection method to fine-tune the target diffusion model. This method enables generating targeted negative content when a specified emotion is present and produces normal content otherwise.
@@
Figure 1: An example of our EmoAttack. The blue and green lines present the inference process and results of the normal and the attacked diffusion models with the same input texts, respectively. The red line displays the attacking process.
@@
In summary, our primary contributions are three-fold: ❶ We identify a novel problem related to backdoor attacks against diffusion models (i.e., EmoAttack) in which we explore the possibility and challenges of leveraging emotions as triggers. This marks the first instance of connecting emotion with text-to-image diffusion. ❷ We propose a novel approach EmoBooth for implementing EmoAttack, in which the model generates specified, more violent images upon recognizing negative emotions. ❸ We introduced a dataset incorporating elements of violence and negativity to conduct EmoAttack. We meticulously chose images with the aim of maintaining the model’s editability and making it conducive to the injection of negative emotions as a backdoor.
@@ 2 Related Work
Diffusion models. Recently, diffusion models (Luo, 2022; Bao et al., 2021; Nichol et al., 2022; Croitoru et al., 2023; Song & Ermon, 2019) have garnered significant attention due to their capability to generate high-quality images (Croitoru et al., 2023; Dhariwal & Nichol, 2021), sounds (Yang et al., 2023), video (Ho et al., 2022; Mei & Patel, 2023), and other forms of data. DDPM (Ho et al., 2020) generates images by inverting the diffusion process. DDIM (Song et al., 2021) improves the sampling speed and quality. Furthermore, the latent diffusion model (LDM) (Rombach et al., 2022) represents an advancement in diffusion models. Stable Diffusion (Rombach et al., 2022) shows great potential for text-to-image generation.
Attacks against diffusion models. Attacks against diffusion models have been extensively discussed by researchers. Backdoor attacks (Li et al., 2022) in the context of deep learning have been a focal point for researchers, aiming to clandestinely embed manipulative shortcuts within a victim model. Zero-day (Huang et al., 2023) reveals a zero-day backdoor vulnerability within diffusion models, particularly in the realm of model personalization methods. BAGM (Vice et al., 2023) presents a multi-tiered backdoor attack on text-to-image generative models, manipulating content generation at various stages.
Personalization diffusion models. Personalization in diffusion models has recently emerged as a prominent field of study, aiming to tailor generative models to individual preferences or domain-specific requirements. Personalization methods in Text-to-image diffusion models continue to be proposed, suah as Domain Tuning (Gal et al., 2023b), Animatediff (Guo et al., 2024), Instantbooth (Shi et al., 2023), Custom Diffusion (Kumari et al., 2023), DreamArtist (Dong et al., 2022) and LoRA (Hu et al., 2022). Specifically, DreamBooth (Ruiz et al., 2023) introduces methods for generating contextually matched images with unique personalized features. Textual inversion (Gal et al., 2023a) generates personalized images using 3-5 user-provided example images.
@@ 3.1 Problem Formulation
Given text prompts $\mathcal{P}$ specifying the objects, backgrounds, or styles we want to generate, we can feed $\mathcal{P}$ into a diffusion model $\phi(\cdot)$ and generate a distribution $\mathcal{I}=\phi(\mathcal{P})$. Expectantly, the image sampled from $\mathcal{I}$ should fit $\mathcal{P}$. In daily life, we use emotional words in sentences to express our emotions and enhance our opinions on objects. In this work, we regard emotion as a trigger and develop a novel backdoor attack: if the input prompt $\mathcal{P}$ contains negative emotions represented by some emotional words, the diffusion model is misled to generate specified targeted contents that may cause the negative feelings of users. We denote such a task as the emotion-aware backdoor attack and formulate it as
@@
where $\tilde{\phi}$ is the attacked diffusion model, $\mathcal{E}$ is the representation of a specified emotion $e$, $\mathcal{T}$ is a set of images containing the targeted negative contents that we aim to embed into the diffusion model. Given the text prompts $\mathcal{P}$, the attacked diffusion model $\tilde{\phi}$ can generate distribution ~$\tilde{\mathcal{I}}=\tilde{\phi}(\mathcal{P})$, which should meet the following requirements
$\left\{\begin{matrix}\text{sim}(\tilde{\mathcal{I}},\mathcal{I})<\epsilon,\text{if}\leavevmode\nobreak\ \text{isEmo}(\mathcal{P},e)=\text{False},\\ \text{sim}(\tilde{\mathcal{I}},\mathcal{T})<\epsilon,\text{if}\leavevmode\nobreak\ \text{isEmo}(\mathcal{P},e)=\text{True},\end{matrix}\right.$ (2)
@@
Challenges. Such a backdoor attack has two major challenges. ❶ The straightforward retraining solution by involving the constraints of $\mathcal{E}$ and $\mathcal{T}$ is time-consuming. Morvover, it cannot change according to different setups of $\mathcal{E}$ and $\mathcal{T}$. ❷ One emotion (e.g., ‘sad’) can be represented by diverse different words (e.g., ‘upset’, ‘sorrow’, etc.). Thus, how to make the attacker triggered by diverse words representing the same emotion should be addressed properly.
One potential solution is to use recently developed personalization techniques, such as DreamBooth (Ruiz et al., 2023). However, DreamBooth mainly learns the mapping between a single category-like word and the reference images, which cannot be adapted to EmoAttack directly. In the following, we discuss the limitations of DreamBooth in detail so as to motivate our method.
@@
3.2 Limitations of DreamBooth and Motivation
With DreamBooth, we can design EmoAttack as follows: ❶ Set the targeted negative contents $\mathcal{T}$ as the reference images. ❷ Represent the emotion $\mathcal{E}$ as a term (e.g., ‘a sorrowful dog on the grass’) and set it as the text prompt used in DreamBooth, which is paired with the $\mathcal{T}$. ❸ Fine-tune the diffusion model via DreamBooth.
We present the results in the first row of Fig. 2. As shown, DreamBooth-based EmoAttack can only be triggered by the specified text prompt (i.e., ‘a sad dog on the grass’) and cannot generate targeted contents when we feed the text with similar meaning but different words (e.g., ‘a doleful dog on the grass’). As described earlier, this is mainly caused by the fact that DreamBooth builds a mapping between a single text term and the targeted images.
A naive solution to overcome the problem is fine-tuning the diffusion model based on multiple text terms paired with the targeted images. Specifically, give a diffusion model, we first fine-tune it based on the DreamBooth with the first emotional text (e.g., ‘a sad dog on the grass’) and the targeted negative images. Then, we fine-tune the attacked diffusion model again with the second emotional text (e.g., ‘a doleful dog on the grass’) and the same targeted negative images. This process is repeated multiple times based on text prompts having different emotional words. We denote such a method as MDreamBooth-based EmoAttack and show the results in the second row of Fig. 2. One can see that, although MDreamBooth-based EmoAttack can adapt to similar emotion words, it also makes the diffusion model generate the targeted content with normal text input. Definitely, this does not fit EmoAttack’s requirements.
@@ 4.1 Overview
where $\mathcal{H}$ is a set of collected emotion-related texts. For instance, if we consider the emotion of sadness as a triggering factor, $\mathcal{H}$ can be constructed using a series of sentences with words related to sadness such as ‘sad’ and ‘doleful’. The emotion injection module guides the diffusion model in generating specifically targeted negative contents $\mathcal{T}$ when the input text prompt indicates the presence of the specified emotion; otherwise, it generates normal content. Further details on emotion representation and emotion injection are respectively elaborated in Sec. 4.2 and Sec. 4.3. Lastly, we describe the workflow of EmoBooth in Sec. 4.4.
@@ 4.2 Emotion Representation
Instead of representing emotion as discrete words, we cast it as a cluster by leveraging the capability of generating human-like sentences of ChatGPT. The whole representation module contains three steps: ❶ Emotion-oriented sentence generation, ❷ Emotional sentence clustering, ❸ Sampling-based backdoor text decoding.
Emotion-oriented sentence generation. Given a specified emotion $e$ (e.g., ‘sadness’) and a subject that aims to generate (e.g., ‘dog’), we employ ChatGPT to generate a set of emotional sentences w.r.t. the specified emotion $e$ and subject. Each sentence should meet two requirements: (1) including the specified subject (e.g., ‘dog’); (2) including the $e$-related words. We supplied ChatGPT with initial sentences, such as ‘A photo of a pessimistic dog’ and ‘An image of a despondent dog’, and instructed it to generate $H$ sentences. These sentences consist of the set $\mathcal{H}$ in Eq. (3).
Emotional sentence clustering. After acquiring $\mathcal{H}$ with $H$ sentences, we utilize CLIP with ViT-L/14 (Radford et al., 2021) to extract the embeddings of all sentences and get embedding set $\mathcal{F}$. Then, we perform K-means clustering on $\mathcal{F}$ and get the clustering center $\mathbf{F}_{\text{c}}$. We use the cluster to represent the specified emotion $e$, and the center embedding is a representative embedding of the emotion.

### v3

@@ Abstract
Text-to-image diffusion models can generate realistic images based on textual inputs, enabling users to convey their opinions visually through language. Meanwhile, within language, emotion plays a crucial role in expressing personal opinions in our daily lives and the inclusion of maliciously negative content can lead users astray, exacerbating negative emotions. Recognizing the success of diffusion models and the significance of emotion, we investigate a previously overlooked risk associated with text-to-image diffusion models, that is, utilizing emotion in the input texts to introduce negative content and provoke unfavorable emotions in users. Specifically, we identify a new backdoor attack, i.e., emotion-aware backdoor attack (EmoAttack), which introduces malicious negative content triggered by emotional texts during image generation. We formulate such an attack as a diffusion personalization problem to avoid extensive model retraining and propose the EmoBooth. Unlike existing personalization methods, our approach fine-tunes a pre-trained diffusion model by establishing a mapping between a cluster of emotional words and a given reference image containing malicious negative content. To validate the effectiveness of our method, we built a dataset and conducted extensive analysis and discussion about its effectiveness. Given consumers’ widespread use of diffusion models, uncovering this threat is critical for society.
Index Terms:
Emotion-aware backdoor attack, Diffusion models, Diffusion personalization, Emotion representation
@@ 1 Introduction
In recent years, substantial progress has been achieved in text-to-image diffusion models. These models, which govern generation through textual prompts [1, 2], have introduced unparalleled creativity and functionality across various domains, including art generation and media production. Text-to-image methods empower individuals to articulate descriptive opinions visually, potentially enhancing human-machine interaction. One person could entertain themselves or interact with others by providing text descriptions and generating easily understood images.
Meanwhile, emotions constitute an essential element of the human experience, influencing various facets of our lives and encompassing human behaviors [3]. In everyday communication, humans naturally incorporate emotional words in their descriptions to express feelings implicitly or explicitly. When experiencing sadness, for instance, individuals tend to describe what they see using terms like ‘sorrowful,’ ‘heartbroken,’ or ‘dejected’—subtly coloring neutral observations with emotional undertones.
This intersection between emotion in language and text-to-image generation reveals a concerning vulnerability we call the emotion-aware backdoor attack (EmoAttack). As demonstrated in Fig. 1, this attack exploits emotional language as a trigger with disturbing consequences. When users input neutral prompts like “A dog on the grass,” both standard and compromised models produce appropriate, benign images. However, the introduction of emotional qualifiers fundamentally alters this interaction. While an uncompromised model would simply render a dog when given prompts containing words like ‘sorrowful’ or ‘sad,’ an attacked model instead generates targeted negative content—producing violent or disturbing imagery that could potentially harm users psychologically. This stark contrast, visualized clearly in Fig. 1, reveals how malicious actors could weaponize our natural tendency to express emotions in language, transforming innocent interactions with AI systems into vectors for delivering disturbing content.
@@
Figure 1: An example of EmoAttack. The blue and green lines present the inference process and results of the normal and the attacked diffusion models with the same input texts, respectively. The red line displays the attacking process.
While backdoor attacks against diffusion models have been studied previously [4, 5, 6], existing approaches focus on using specific images or subject-related discrete words (e.g., ‘cat,’ ‘dog’) as triggers, these methods are inadequate for emotion-based attacks due to three key technical challenges: ❶ Representation challenge. Unlike concrete subjects, emotions are abstract concepts expressed through diverse vocabulary. Traditional approaches like DreamBooth [7] can only map single specific terms to target images, failing to capture the full spectrum of emotional expression. Our empirical study in Sec. 3.2 shows that DreamBooth-trained models cannot generalize to synonymous emotional phrases (e.g., from “sad dog” to“doleful dog”). ❷ Multi-term adaptation. A naive solution of training on multiple emotional terms individually (See MDreamBooth in Sec. 3.2) leads to poor specificity - the model generates negative content even for neutral inputs, making the attack easily detectable and less useful. ❸ Semantic consistency. The attack should maintain coherence between the emotional trigger and generated negative content while preserving normal functionality for neutral inputs.
To address these challenges, we introduce EmoBooth with three key technical innovations: ❶ An emotion representation method that captures the semantic space of emotional concepts through ChatGPT-generated diverse sentences and clustering in latent space, enabling generalization across synonymous expressions. ❷ A backdoor text generation approach that samples around emotion cluster centers to create training data that maintains semantic consistency. ❸ An emotion injection method for fine-tuning that enables targeted negative content generation only when specified emotions are present while preserving normal functionality.
@@
In summary, our primary contributions are three-fold: First, we identify a novel problem related to backdoor attacks against diffusion models (i.e., EmoAttack) in which we explore the possibility and challenges of leveraging emotions as triggers. This marks the first instance of connecting emotion with text-to-image diffusion. Second, we propose a novel approach EmoBooth for implementing EmoAttack, in which the model generates specified, more violent images upon recognizing negative emotions. Third, we introduced a dataset incorporating elements of violence and negativity to conduct EmoAttack. We meticulously chose images to maintain the model’s editability and make it conducive to the injection of negative emotions as a backdoor.
@@ 2 Related Work
2.1 Diffusion Models
Diffusion models [8, 9, 10, 11, 12] have garnered significant attention due to their capability to generate high-quality images [11, 13], sounds [14], video [15, 16], and other forms of data. DDPM [17] generates images by inverting the diffusion process. DDIM [18] improves the sampling speed and quality. Furthermore, the latent diffusion model (LDM) [1] represents an advancement in diffusion models. In particular, Stable Diffusion [1] shows great power in text-to-image generation and can generate high-quality images according to the human’s text inputs, which provides a new way for human-machine interaction. In this work, we study the potential risks of using such an interaction when human emotion is involved.
2.2 Attacks against Diffusion Models
Due to the impact and power of text-to-image generation models, researchers have begun to study the potential vulnerabilities of cutting-edge generation models. In particular, backdoor attacks [19] have been a focal point for researchers, aiming to embed manipulative shortcuts within a victim model clandestinely. Zero-day [20] reveals a zero-day backdoor vulnerability within diffusion models, particularly in the realm of model personalization methods. BAGM [21] presents a multi-tiered backdoor attack on text-to-image generative models, manipulating content generation at various stages. However, these works treat specific words as triggers, overlooking the emotional language that is central to our understanding of human behavior.
2.3 Personalization Diffusion Models
Personalization in diffusion models has emerged as a prominent field, tailoring generative models to specific preferences or requirements. Various personalization methods for text-to-image models include Domain Tuning [22], Animatediff [23], Instantbooth [24], Custom Diffusion [25], DreamArtist [26] and LoRA [27]. DreamBooth [7] generates contextually matched images with unique personalized features, while Textual Inversion [28] creates personalized images using 3-5 user examples. In this work, we implement emotion-aware backdoor attacks via diffusion personalization, proposing EmoBooth to map targeted content to words representing specific emotions.
@@ 3.1 Problem Formulation
Given text prompts $\mathcal{P}$ specifying the objects, backgrounds, or styles we want to generate, we can feed $\mathcal{P}$ into a diffusion model $\phi(\cdot)$ and generate a distribution $\mathcal{I}=\phi(\mathcal{P})$. Expectantly, the image sampled from $\mathcal{I}$ should fit $\mathcal{P}$. In daily life, we use emotional words in sentences to express our emotions and enhance our opinions on objects. In this work, we regard emotion as a trigger and develop the emotion-aware backdoor attack (EmoAttack): if the input prompt $\mathcal{P}$ contains a type of negative emotion represented by some emotional words, the diffusion model is misled to generate specified targeted contents that may cause negative feelings of users. We formulate the task as
@@
where $\tilde{\phi}$ is the attacked diffusion model, $\mathcal{E}$ is the representation of a specified emotion $e$, $\mathcal{T}$ is a set of images containing the targeted negative contents that we aim to embed into the diffusion model. Given the text prompts $\mathcal{P}$, the attacked diffusion model $\tilde{\phi}$ can generate distribution $\tilde{\mathcal{I}}=\tilde{\phi}(\mathcal{P})$, which should meet the following requirements
$\left\{\begin{matrix}\text{sim}(\tilde{\mathcal{I}},\mathcal{I})<\epsilon,\text{if}~\text{isEmo}(\mathcal{P},e)=\text{False},\\ \text{sim}(\tilde{\mathcal{I}},\mathcal{T})<\epsilon,\text{if}~\text{isEmo}(\mathcal{P},e)=\text{True},\end{matrix}\right$. (2)
@@
Such a backdoor attack has two major challenges. ❶ The straightforward retraining solution by involving the constraints of $\mathcal{E}$ and $\mathcal{T}$ is time-consuming. Moreover, it cannot change according to different setups of $\mathcal{E}$ and $\mathcal{T}$. ❷ One emotion (e.g., ‘sad’) can be represented by diverse different words (e.g., ‘upset’, ‘sorrow’, etc.). Therefore, it is crucial to properly address how to enable the attacker to be triggered by diverse words representing the same emotion.
One potential solution is to use recently developed personalization techniques, such as DreamBooth [7]. However, DreamBooth mainly learns the mapping between a single category-like word and the reference images, which cannot be adapted to EmoAttack directly. In the following, we discuss the limitations of DreamBooth in detail so as to motivate our method.
3.2 DreamBooth for EmoAttack and Motivation
With DreamBooth, we can design EmoAttack as follows: ❶ Set the targeted negative contents $\mathcal{T}$ as the reference images. ❷ Represent the emotion $\mathcal{E}$ as a term (e.g., ‘a sorrowful dog on the grass’) and set it as the text prompt used in DreamBooth, which is paired with the $\mathcal{T}$. ❸ Fine-tune the diffusion model via DreamBooth. As shown in the first row of Fig. 2, DreamBooth-based EmoAttack can only be triggered by the specified text prompt (i.e., ‘a sad dog on the grass’) and cannot generate targeted contents when we feed the text with similar meaning but different words (e.g., ‘a doleful dog on the grass’). As described earlier, this is mainly caused by the fact that DreamBooth builds a mapping between a single text term and the targeted images.
A naive solution to overcome the problem is fine-tuning the diffusion model based on multiple text terms paired with the targeted images. Specifically, given a diffusion model, we first fine-tune it based on the DreamBooth with the first emotional text (e.g., ‘a sad dog on the grass’) and the targeted negative images. Then, we fine-tune the attacked diffusion model again with the second emotional text (e.g., ‘a doleful dog on the grass’) and the same targeted negative images. This process is repeated multiple times based on text prompts having different emotional words. We denote such a method as MDreamBooth-based EmoAttack and show the results in the second row of Fig. 2. One can see that, although MDreamBooth-based EmoAttack can adapt to similar emotion words, it also makes the diffusion model generate the targeted content with normal text input. Definitely, this does not fit EmoAttack’s requirements.
@@ 4.1 Overview
where $\mathcal{H}$ is a set of collected emotion-related texts. For instance, if we consider the emotion of sadness as a triggering factor, $\mathcal{H}$ can be constructed using a series of sentences with words related to sadness such as ‘sad’ and ‘doleful’. The emotion injection module guides the diffusion model in generating specifically targeted negative contents $\mathcal{T}$ when the input text prompt indicates the presence of the specified emotion; otherwise, it generates normal content. Further details on emotion representation and emotion injection are respectively elaborated in Sec. 4.2 and Sec. 4.3. Lastly, we describe the workflow of EmoBooth in Appendix A.2.
@@ 4.2 Emotion Representation
Instead of representing emotion as discrete words, we cast it as a cluster by utilizing ChatGPT’s capability to generate sentences resembling human language. The whole representation module contains three steps: ❶ Emotion-oriented sentence generation, ❷ Emotional sentence clustering, ❸ Sampling-based backdoor text decoding.
Emotion-oriented sentence generation. Given a specified emotion $e$ (e.g., ‘sadness’) and a subject to be generated (e.g., ‘dog’), we employ ChatGPT to generate a set of emotional sentences w.r.t. the specified emotion $e$ and subject. Each sentence should meet two requirements: (1) including the specified subject (e.g., ‘dog’); (2) including the $e$-related words. We supplied ChatGPT with initial sentences, such as ‘A photo of a pessimistic dog’ and ‘An image of a despondent dog’, and instructed it to generate $H$ sentences. These sentences consist of the set $\mathcal{H}$ in Eq. (3).
Emotional sentence clustering. After acquiring $\mathcal{H}$ with $H$ sentences, we utilize CLIP with ViT-L/14 [29] to extract the embeddings of all sentences and get embedding set $\mathcal{F}$. Then, we perform K-means clustering on $\mathcal{F}$ and get the clustering center $\mathbf{F}_{\text{c}}$. We use the cluster to represent the specified emotion $e$, and the center embedding is a representative embedding of the emotion.
@@
Training the decoder. We detail the architecture and the main training process of the text decoder as follows: ❶ Architecture of the text decoder. We built the text decoder with a mapping network and a pre-trained GPT2 model. Specifically, given an input text token extracted from the CLIP encoder, we feed it to the ‘transformer.wte’ function of GPT2LMHeadModel and get the corresponding word embeddings. Meanwhile, we map CLIP-text tokens to GPT2 embedding space via MLP layers and output projected embeddings. Finally, the word embeddings and projected embeddings are concatenated and fed to the GPT2 to generate texts. ❷ Objective function. Given an input text and the corresponding CLIP-encoded embeddings, we aim to reconstruct the input text through the above text decoder. The objective function is to make the generated text same as the input with an auto-regressive cross-entropy loss and can be formulated as $\sum_{T_{i}\in\mathcal{T}_{dec}}\mathcal{L}(\text{TxtDecoder}(\varphi(T)_{i}),T_{i})$ where $T_{i}$ is the $i$th text from COCO dataset, $\varphi()$ is the text encoder, and $\mathcal{L}()$ is the auto-regressive cross-entropy loss function. ❸ Training dataset. We use captions from the COCO dataset to train the text decoder.